a) A beam of 7.8-MeV  particles scatters from a gold foil of thickness 0.30 m.
i) What fraction of the  particles is scattered between 1.5° and 2.5°?
ii) What is the ratio of  particles scattered through angles greater than 2° to the number
scattered through angles greater than 10°?
1
Expert's answer
2020-06-15T10:21:43-0400

Given:

"Z_1=2,\\
Z_2=79,
\\t=30\cdot10^{-6}\text{ m},\\
\rho=19300\text{ km/m}^3,\\
\Mu=197\cdot10^{-3}\text{ kg/mol},\\
\theta_1=1.5°,\\
\theta_2=2.5°,\\
\theta_3=2°,\\
\theta_4=10°."


Solution

The fraction of particles scattered at angle θ is


"f(\theta)=\pi nt\bigg(\frac{Z_1Z_2e^2}{8\pi\epsilon_0K}\bigg)^2\text{cot}^2\bigg(\frac{\theta}{2}\bigg)."


In this equation, t - thickness, K - kinetic energy, n is the number of atoms per unit volume of material (gold):


"n=\frac{\rho N_A}{\Mu}=5.9\cdot10^{28}\text{ atoms/m}^3,"

where "\Mu" is the molar mass of Gold and "\rho" is its density.


i) Between 1.5° and 2.5° the fraction will be


"f(\theta_1)-f(\theta_2)=\pi nt\bigg(\frac{Z_1Z_2e^2}{8\pi\epsilon_0K}\bigg)^2\bigg[\text{cot}^2\bigg(\frac{\theta_1}{2}\bigg)-\text{cot}^2\bigg(\frac{\theta_2}{2}\bigg)\bigg]=\\\space\\
=0.0442."

ii) Now it is much easier to find the ratio of particles scattered through angles greater than 2° to the number scattered through angles greater than 10°:


"\frac{f(\theta_3)}{f(\theta_4)}=\frac{\text{cot}^2\frac{\theta_3}{2}}{\text{cot}^2\frac{\theta_4}{2}}=25."